# Mixture of Contexts for Long Video Generation

### v2

@@ 2 Related Work
Long Video Generation. Existing video generation models [23, 37, 13, 49, 15, 1, 17, 5, 6, 12, 53] are mostly limited to a few seconds. Several recent frameworks specifically target longer video generation using autoregressive models that operate on frames, chunks, or segments, such as CausVid [50]. While these frameworks extend generation capabilities, they often grapple with error accumulation [39] inherent in sequential prediction or face uncertain computational scaling to longer durations. To mitigate these issues, RollingDiffusion [32] and Diffusion Forcing [3] inject controlled noise into the historical context and train the model to denoise it, increasing robustness to compounding errors. MAGI-1 [33] and SkyReels-V2 [4] scale up these ideas by employing autoregressive denoising, aiming for potentially longer durations. An orthogonal strategy is to distill the entire past into a constant-size latent. TTTVideo [7] and LaCT[60] use a learnable MLP to encode the context during inference, while FramePack [57] encodes arbitrarily many frames into a fixed vector for next-frame prediction. FramePack [57] also proposes early planning of future frames to mitigate the error accumulation issue. This is similar to using keyframes or anchor frames [16, 40, 27, 62, 19, 47, 48, 45], where certain frames are predefined and the video generation model only does an interpolation sampling job. These methods extend video generation to the one-minute range but still face a hard ceiling on maintaining long-context coherence going forward, as they rely on lossy compression of the contexts. The work most closely related to ours is Long-Context Tuning [14] (LCT), which starts from a single-shot DiT and expands its context window to a scene comprising up to eight shots ($\approx$8s, $\sim 2.3{\times}10^{4}$ tokens each). LCT [14] keeps the attention mechanism dense: all text and video tokens inside the enlarged window attend to one another after being positioned with an interleaved 3D RoPE. While this design elegantly re-uses the pretrained weights and yields impressive multi-shot coherence, it inherits the quadratic cost of full self-attention – FLOPs and memory scale with $(8L_{\text{shot}})^{2}$.
@@ 3.1 Mixture of Contexts
where $\cdot$] concatenate and join all indices of the top-$k$ chunks. The relevance between the $\bm{q}_{i}$ and the chunk sequence $\bm{K}_{\omega}$ is determined by the inner product of $\bm{q}_{i}$ and the descriptor for $\bm{K}_{\omega}$ denoted as $\phi(\bm{K}_{\omega})$. For this work, we use the simple, efficient, yet effective mean pooling operation as the descriptor transformation $\phi$. We argue that such a mean pooling operation is highly sufficient and expressive for video generation tasks. First, mean pooling is well-motivated by established practices in multimodal matching, such as in CLIP [31], where global image representations are derived by averaging token embeddings to compute similarity with text embeddings. This approach effectively captures dominant semantic features while being robust to local variations, a property that translates naturally to video chunks where spatially and temporally adjacent tokens often represent redundant or correlated visual elements (e.g., static backgrounds or gradual motions). In our trainable framework, mean pooling is not a static heuristic but an adaptive mechanism: while top-$k$ itself is non-differentiable, the model learns indirectly through the attention mechanism on selected chunks. Specifically, if a selected chunk proves irrelevant during attention computation, gradients from the loss will flow back through its keys/values, which is the source of the mean-pooled descriptor. This process attenuates unhelpful representations and encourages the query/key projections to produce more discriminative similarities over training iterations. This self-correcting process aligns with indirect adaptation seen in hard-routing MoE systems and sparse attention frameworks (e.g., where downstream modules provide the learning signal despite discrete and non-differentiable selections). This end-to-end differentiability and parameter-less router ensures that the seemingly simple dot-product routing becomes highly expressive, as the network shapes embeddings to emphasize discriminative features for sparse attention, without introducing additional parameters or computational overhead. Empirical zero-shot application to pretrained models further validates its efficacy, as will be detailed in our supplementary material.
@@
Context Drop-off and Drop-in. To enhance the robustness of our Mixture of Contexts (MoC) and mitigate issues akin to the “dead expert” problem in Mixture-of-Experts (MoE) systems, where certain experts or routes may become underutilized or ignored during training, we introduce two complementary regularization techniques: context drop off and context drop in. These perturbations are applied stochastically during training to simulate noisy or imperfect routing decisions, encouraging the model to learn more resilient representations that do not over-rely on specific chunks while ensuring balanced utilization across the context space. Motivated by the observation that routing may suffer from inaccuracies due to noise in embeddings or evolving data distributions, context drop off randomly removes a subset of the top-$k$ selected chunks for each query token. Specifically, for a given query $q_{i}$, after computing the routed indices $\Omega(q_{i})$ in Eq. 3, we sample a drop probability $p_{\text{drop}}\sim\text{Uniform}(0,p_{\max})$ and mask out $\lfloor p_{\text{drop}}\cdot k\rfloor$ randomly chosen chunks from $\Omega(q_{i})$. This forces the model to generate coherent outputs even when certain chosen context is sporadically unavailable, promoting redundancy in the learned dependencies and preventing catastrophic failure from routing errors. Conversely, context drop in injects extraneous chunks into the selected set to simulate over-inclusive routing. For each query, we randomly sample $m\sim\text{Poisson}(\lambda)$ chunks to be included in the selected pool $\Omega(q_{i})$. This technique combats the dead route problem by artificially activating underutilized chunks, ensuring gradients flow through a broader range of context segments and balancing the routing distribution over time. Since our router is parameter-less and relies solely on mean-pooled feature similarity, these regularization techniques do not interfere with the learning of the routing mechanism itself. If a chunk is truly important, its relevance will be naturally enhanced through backpropagation in the attention modules, as the model adjusts the query and key projections to amplify meaningful similarities and de-emphasize irrelevant ones. In essence, the end-to-end differentiability of the system means that the attention process implicitly serves as the router’s learning signal, making the framework self-correcting and adaptive without dedicated routing parameters.
@@ 4 Experiments
We argue that such a mean pooling operation is highly suitable for videos since pixels that lie close in space and neighboring frames tend to depict the same object or background region. After the DiT [29]’s patch embedding, these tokens occupy a very narrow subspace: their first principal component often explains >90% of the local variance in practice. The arithmetic mean is exactly that first-component estimator for centered data, so a simple average already captures the dominant semantics of the whole chunk while discarding high-frequency noise. Zero-shot experiments support this claim – applying such a routing strategy directly to a pretrained video generation model, as will be shown in our supplementary material. Although the routing score in Eq. 3 is literally just a dot-product between the query and a mean-pooled key, it is not a fixed heuristic: the key vectors being averaged and the query vector doing the scoring are both produced by weights that are updated during training. Gradients flow through the mean-pool operation and the subsequent top-$k$ mask back to the projection matrices, allowing the model to learn how to shape each chunk’s pooled key and each query in a way that best separates useful from irrelevant context. In practice, this makes the ostensibly “simple” mean + top-$k$ rule highly expressive without introducing extra routing parameters or computation, as the network continuously adapts its internal representations to exploit it.
MoC Implementation Benchmark.
Figure 4: Performance benchmark of our content-aligned Mixture of Contexts implementation with full attention (implemented with Flash Attention 2 [9, 8]). Our method stays near linear with respect to the shot number (xaxis, assuming 8 seconds, 12 FPS, roughly 23k tokens), or in other words, the sequence length $L$.
We benchmark our adaptive MoC’s performance with full attention (implemented with Flash Attention 2 [9, 8]) in Fig. 4, where our method stays near-linear in terms of FLOPs and latency with respect to the number of shots, or in other words, the sequence length $L$. On top of sparsity, the key to this efficiency lies in three design decisions: (1) the use of on-the-fly segment_reduce pooling avoids materializing variable-length chunks in memory; (2) tokens are organized in head-major order to ensure coalesced memory access during gather operations; and (3) the entire routing + attention computation is wrapped in a single Flash Attention [9, 8] var-len call, preserving kernel fusion and minimizing overhead.

### v3

@@ 2 Related Work
Long Video Generation. Existing video generation models [23, 36, 13, 50, 15, 1, 17, 5, 6, 12, 56] are mostly limited to a few seconds. To push beyond this short horizon, TECO [48] introduces temporally consistent transformers with a recurrent state to propagate information over long sequences, and NUWA-XL [51] adopts a diffusion-over-diffusion hierarchy that generates extremely long videos by first synthesizing sparse keyframes and then recursively filling in between. Several recent frameworks specifically target longer video generation using autoregressive models that operate on frames, chunks, or segments, such as MALT [54] and CausVid [52]. While these frameworks extend generation capabilities, they often grapple with error accumulation [38] inherent in sequential prediction or face uncertain computational scaling to longer durations. To mitigate these issues, RollingDiffusion [31] and Diffusion Forcing [3] inject controlled noise into the historical context and train the model to denoise it, increasing robustness to compounding errors. MAGI-1 [32] and SkyReels-V2 [4] scale up these ideas by employing autoregressive denoising, aiming for potentially longer durations. An orthogonal strategy is to distill the entire past into a constant-size latent. TTTVideo [7] and LaCT[63] use a learnable MLP to encode the context during inference, while FramePack [60] encodes arbitrarily many frames into a fixed vector for next-frame prediction. FramePack [60] also proposes early planning of future frames to mitigate the error accumulation issue. This is similar to using keyframes or anchor frames [16, 39, 27, 65, 19, 47, 49, 45], where certain frames are predefined and the video generation model only does an interpolation sampling job. These methods extend video generation to the one-minute range but still face a hard ceiling on maintaining long-context coherence going forward, as they rely on lossy compression of the contexts. The work most closely related to ours is Long-Context Tuning [14] (LCT), which starts from a single-shot DiT and expands its context window to a scene comprising up to eight shots ($\approx$8s, $\sim 2.3{\times}10^{4}$ tokens each). LCT [14] keeps the attention mechanism dense: all text and video tokens inside the enlarged window attend to one another after being positioned with an interleaved 3D RoPE. While this design elegantly re-uses the pretrained weights and yields impressive multi-shot coherence, it inherits the quadratic cost of full self-attention – FLOPs and memory scale with $(8L_{\text{shot}})^{2}$.
@@ 3.1 Mixture of Contexts
where $\cdot$] concatenate and join all indices of the top-$k$ chunks. The relevance between the $\bm{q}_{i}$ and the chunk sequence $\bm{K}_{\omega}$ is determined by the inner product of $\bm{q}_{i}$ and the descriptor for $\bm{K}_{\omega}$ denoted as $\phi(\bm{K}_{\omega})$. For this work, we use the simple, efficient, yet effective mean pooling operation as the descriptor transformation $\phi$. We argue that such a mean pooling operation is highly sufficient and expressive for video generation tasks. The effectiveness of this design relies on the intrinsic quality of the Diffusion Transformer’s learned features. As demonstrated by DDAE [43], denoising diffusion autoencoders function as unified self-supervised learners, naturally acquiring semantically meaningful and linearly separable internal representations. Consequently, the global average of a token chunk effectively captures its dominant semantic content and visual layout, providing a robust summary that enables queries to distinguish relevant context based on high-level alignment. This approach effectively captures dominant semantic features while being robust to local variations, a property that translates naturally to video chunks where spatially and temporally adjacent tokens often represent redundant or correlated visual elements (e.g., static backgrounds or gradual motions). Furthermore, in our trainable framework, this pooling is not a static heuristic but an adaptive mechanism: while top-$k$ itself is non-differentiable, the model learns indirectly through the attention mechanism on selected chunks. Specifically, if a selected chunk proves irrelevant during attention computation, gradients from the loss will flow back through its keys/values, which is the source of the mean-pooled descriptor. This process attenuates unhelpful representations and encourages the query/key projections to produce more discriminative similarities over training iterations. This self-correcting process aligns with indirect adaptation seen in hard-routing MoE systems and sparse attention frameworks (e.g., where downstream modules provide the learning signal despite discrete and non-differentiable selections). This end-to-end differentiability and parameter-less router ensures that the seemingly simple dot-product routing becomes highly expressive, as the network shapes embeddings to emphasize discriminative features for sparse attention, without introducing additional parameters or computational overhead. Empirical zero-shot application to pretrained models further validates its efficacy, as will be detailed in our supplementary material.
@@
Context Drop-off. To enhance the robustness of our Mixture of Contexts (MoC) and mitigate issues akin to the “dead expert” problem in Mixture-of-Experts (MoE) systems, we first introduce context drop-off. Motivated by the observation that routing may suffer from inaccuracies due to noise in embeddings or evolving data distributions, this technique randomly removes a subset of the top-$k$ selected chunks for each query token. Specifically, for a given query $q_{i}$, after computing the routed indices $\Omega(q_{i})$ in Eq. 3, we sample a drop probability $p_{\text{drop}}\sim\text{Uniform}(0,p_{\max})$ and mask out $\lfloor p_{\text{drop}}\cdot k\rfloor$ randomly chosen chunks from $\Omega(q_{i})$. This forces the model to generate coherent outputs even when a certain chosen context is sporadically unavailable, promoting redundancy in the learned dependencies and preventing catastrophic failure from routing errors.
Context Drop-in. Complementarily, we employ context drop-in to inject extraneous chunks into the selected set to simulate over-inclusive routing. For each query, we randomly sample $m\sim\text{Poisson}(\lambda)$ chunks to be included in the selected pool $\Omega(q_{i})$. This technique combats the dead route problem by artificially activating underutilized chunks, ensuring gradients flow through a broader range of context segments and balancing the routing distribution over time. Since our router is parameter-less and relies solely on mean-pooled feature similarity, these regularization techniques do not interfere with the learning of the routing mechanism itself. Instead, if a chunk is truly important, its relevance will be naturally enhanced through backpropagation in the attention modules, as the model adjusts the query and key projections to amplify meaningful similarities. In essence, the end-to-end differentiability of the system means that the attention process implicitly serves as the router’s learning signal, making the framework self-correcting and adaptive without dedicated routing parameters.
Per-Head Distributed Routing. A crucial design choice in Mixture of Contexts is the granularity of the retrieval process: is context selected globally once, or dynamically at every step? We implement routing at the finest granularity – independently for each attention head in every layer. Rather than relying on a single “global” router to select a fixed set of $k$ chunks shared across the entire network, our approach effectively acts as an ensemble of $L_{\text{layers}}\times H_{\text{heads}}$ independent routers. This distinction is vital for two reasons. First, different attention heads in diffusion transformers specialize in distinct feature subspaces (e.g., low-level texture coherence vs. high-level semantic identity), necessitating access to different historical segments. Second, while each head is strictly sparse (attending to only $k$ chunks), the union of selected chunks across all heads and layers covers a significantly larger portion of the context. This distributed routing ensures sufficient global communication and prevents the information bottleneck that would arise from a static global selection, allowing the model to utilize its entire parameter space to reconstruct the full context manifold through diverse, sparse viewpoints.
@@ 4 Experiments
Figure 4: Illustration of coherence achieved by Mixture of Contexts. We highlight specific visual elements (indicated by red and green circles) that persist faithfully across shots, demonstrating the efficacy of MoC’s retrieval mechanism. Row 1: Background landmarks (cityscape buildings) remain geometrically consistent despite camera movement. Row 2: Semantic consistency is maintained where a sketchbook drawing transitions into the corresponding physical architecture. Row 3: Spatial layout and background elements (neon signage) are preserved across reverse-angle cuts. Row 4: Fine-grained object identity is retained, such as the side vent structure (green) and screen (red) of the computer. Row 5: Multi-character consistency is achieved in a dynamic car interior, where the identities of the driver, passenger, and child remain distinct without feature mixing.
Qualitative Results. We present qualitative comparisons in Fig. 3. We argue that such a mean pooling operation is highly suitable for videos since pixels that lie close in space and neighboring frames tend to depict the same object or background region. After the DiT [29]’s patch embedding, these tokens occupy a very narrow subspace: their first principal component often explains >90% of the local variance in practice. The arithmetic mean is exactly that first-component estimator for centered data, so a simple average already captures the dominant semantics of the whole chunk while discarding high-frequency noise. Zero-shot experiments support this claim – applying such a routing strategy directly to a pretrained video generation model, as will be shown in our supplementary material. Although the routing score in Eq. 3 is literally just a dot-product between the query and a mean-pooled key, it is not a fixed heuristic: the key vectors being averaged and the query vector doing the scoring are both produced by weights that are updated during training. Gradients flow through the mean-pool operation and the subsequent top-$k$ mask back to the projection matrices, allowing the model to learn how to shape each chunk’s pooled key and each query in a way that best separates useful from irrelevant context. In practice, this makes the ostensibly “simple” mean + top-$k$ rule highly expressive without introducing extra routing parameters or computation, as the network continuously adapts its internal representations to exploit it.
Qualitative Illustration of Coherence. We provide a qualitative verification of long-term coherence in Fig. 4, demonstrating that Mixture of Contexts (MoC) robustly preserves consistency across diverse modalities and shot boundaries. The visualizations confirm that our learned dynamic sparse attention routing mechanism effectively maintains geometric background stability, fine-grained object details, and semantic alignment throughout the generation process. Furthermore, the model demonstrates strong multi-character/subject consistency, successfully distinguishing and preserving the unique identities of multiple subjects in dynamic scenes without feature mixing or identity drifting. MoC successfully retrieved these small details and highly abstracted semantic contents across hundreds even thousands of frames.